The probability is 0.45 that a traffic fatality involves an intoxicated or​ alcohol-impaired driver or nonoccupant. In six traffic​ fatalities, find the probability that the​ number, Y, which involve an intoxicated or​ alcohol-impaired driver or nonoccupant is a. exactly​ three; at least​ three; at most three. b. between two and four​, inclusive. c. Find and interpret the mean of the random variable Y. d. Obtain the standard deviation of Y.
Answer #1

Solution:- Given that p =0.45, n = 6

a.

P(Y = 3) = 6C3*0.45^3*0.55^3 = 0.3032

P(Y >= 3) = 0.5585

P(Y <= 3) = 0.7447

b. P(2 <= x <= 4) = P(2) + P(3) + P(4) = 0.2780+0.3032+0.1861 = 0.7673

c. mean = n*p = 6*0.45 = 2.7

d. standard deviation = sqrt(npq) = sqrt(6*0.45*0.55) = 1.2186